# WARP-LCA: Efficient Convolutional Sparse Coding with Locally Competitive Algorithm

### v1

@@ 4 Experiments and Results
In our experiments we used the CIFAR-10 dataset (Krizhevsky et al., 2009) for learning both the dictionary and the encoder model (WARP-CNN). The dictionary was learned using a soft thresholding LCA approach (you may find the dictionary here Figure 6). Additionally, to test the versatility of our method on out-of-distribution data we employed the Oxford Pets dataset (Parkhi et al., 2012). All of our experiments were conducted on a Linux system equipped with three NVIDIA Tesla V100-PCIE-16GB GPUs, each with a compute capability of 7.0 and 16.94 GB of memory. The system also features 8 CPU cores and memory resources, with a total of 201.18 GB of RAM and 174.20 GB available. On average, training took 15.6 iterations per second using batches of 32 images (for dictionary training on the CIFAR dataset, with 200 LCA iterations).
@@
Performance was assessed using Mean Squared Error (MSE), $\ell_{0}$-norm, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM). These metrics were chosen to provide a comprehensive view of both quantitative and qualitative measures. Each model was executed for 1,000 iterations with lambdas set at 0.15 and 0.85, alongside several intermediate levels to analyze relative speed improvements based on maximum PSNR achieved by the LCA after 1000 iterations. Execution of the models, i.e. inference allows us to change the number of iterations. To compare both WARP-LCA and LCA up to full convergence, we have evaluated inference after 1000 iterations.
@@
Figure 2: Encoding Performance Comparison between WARP-LCA and LCA at $\lambda=0.2$. The figure displays the performance metrics for WARP-LCA and LCA. The columns represent different metrics: Log Mean Squared Error (MSE), $\ell_{0}$-Norm, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM). For each plot, WARP-LCA and LCA Mean and Standard Deviation of the Metrics are shown for comparison.
@@
Table 1: Evaluation after 1000 iterations on CIFAR-10 dataset with $\lambda=0.15$
Metric WARP-LCA Metrics Mean LCA Metrics Mean
Recon. MSE 0.0108 0.0129
$L_{0}$ Norm 864.62 924.81
$L_{1}$ Norm 466.86 422.34
PSNR 30.92 30.12
SSIM 0.9467 0.9398
@@
Our results, presented in Figure 2 and in Table 1, highlight significant findings. The WARP-LCA demonstrated markedly faster convergence in terms of MSE as indicated by the log(MSE) trajectory, maintaining a superior MSE even after 1,000 iterations compared to the standard LCA. Similarly, $\ell_{0}$-norm results showed that WARP-LCA began near optimal sparsity levels, whereas LCA required considerable iterations to approximate these levels. Markedly the $\ell_{0}$-norm achieved by the WARP-LCA is lower than $\ell_{0}$-norm reached by the LCA. Both PSNR and SSIM metrics reflected faster convergence and higher final values for the WARP-LCA.
Figure 3: Relative Speed Up of WARP-LCA Compared to LCA. The graph quantifies the efficiency of WARP-LCA relative to LCA in achieving equivalent PSNR levels, using the CIFAR-10 test set for evaluation. The comparison point for PSNR is established after 1000 iterations of the LCA. The vertical axis depicts the reduction in the number of iterations required by WARP-LCA to match the PSNR achieved by LCA.
@@ 4.1 Denoising Performance in Classification Pipelines
To evaluate the robustness of our WARP-LCA model in enhancing the denoising capabilities within classification pipelines, we employed three distinct pre-trained convolutional neural network architectures: DenseNet-40 with a growth rate of 12 on CIFAR-10 (Huang et al., 2017), Wide Residual Network with depth 40 and widening factor 8 on CIFAR-10 (Zagoruyko and Komodakis, 2016), and ResNeXt with 29 layers and cardinality 32 using 4d expansions (Xie et al., 2017).
The test set images from CIFAR-10 were subjected to varying levels of additive Gaussian noise. For each noise level and each backbone architecture, two denoising strategies were applied: traditional LCA-based denoising, run for 800 iterations, and our WARP-LCA-based denoising, run for 200 iterations. This approach allowed us to directly compare the efficacy of our WARP-LCA model against the standard LCA under identical noise conditions, consistent with earlier experiments demonstrating comparable PSNR values between the models. Although different levels of noise typically require adjustments to the sparsity parameter $\lambda$, we opted for a $\lambda$ of 0.2 across all noise levels, drawing from methodologies similar to those reported by Teti et al. (2022). Denoising performance was quantitatively assessed by measuring the classification accuracy post-denoising.
@@
As shown in Figure 4, WARP-LCA consistently outperforms traditional LCA in terms of classification accuracy across all tested noise levels and models.
@@ 5 Discussion
The comparative analysis in Section 4 (illustrated in Figure 2) compellingly demonstrates the superior efficacy of WARP-LCA, over the traditional LCA: faster convergence and settling for deeper minima. This attribute was consistently evident across a range of noise levels and architectural backbones in the denoising experiment in Section 4.1. These results not only corroborate but also extend the findings by Teti et al. (2022), propelling WARP-LCA beyond the current state-of-the-art. This advancement underscores the potential of predictive initializations to redefine performance benchmarks in the field, offering robust, fast, and more accurate sparse coding.
The robustness of WARP-LCA, particularly for out-of-distribution data, was tested on the Oxford Pets dataset (Parkhi et al., 2012), which starkly contrasts with CIFAR-10 in terms of image size (images in the Oxford Pets datset are much larger) and image statistics. The Oxford Pets dataset is a collection of images featuring 37 different breeds of cats and dogs, each annotated with class labels, bounding boxes, and pixel-level segmentation masks. Despite these considerable differences, and the fact that our models were initially trained only on CIFAR images, the fully convolutional nature of predictor architectures allowed for an efficient adaptation to this new context.
@@
The (relative) accumulated activation map generated by WARP-LCA provides more homogeneous and less extreme structures than the accumulated LCA activation map, indicated much more uniformly distributed activations. Even after 300 iterations, activity in the accumulated LCA map is still concentrated in very dark and light areas of the image. In addition, WARP-LCA also has a lower $\ell_{0}$-Norm than the LCA. The reconstructed image from the LCA appears not to have converged yet, as the colors, especially in the light and dark areas, seems to be faded. The reconstructed image from the WARP-LCA seems to be much closer to the original. This may imply that the CIFAR dictionary alone may be sufficient to be used on data other than CIFAR images and WARP-LCA managed to predict states that lead to a consistent and balanced activation profile.
We have experimented with different approaches to warm-starting LCA, including random initialization but found that the only methods that lead to convergence are the conventional zero initialization as is typical for LCA and our CNN predictor. We have compared the similarity of learned cifar10-kernels to the correlation of activations of LCA and WARP-LCA after 1000 optimization cycles up to convergence and found that the WARP-LCA method has lower correlation between kernels as would be expected from the earlier results. The prediction of initial states leads to more diverse and robust minima (please see the appendix for a comparrison of LCA, WARP-LCA, CNN-only prediction and simple convolution).
@@
It is important to note that our experiments were conducted using a single, learned over-complete dictionary. The application of WARP-LCA to dictionaries with lower overcompleteness—and hence fewer kernels—might prove more challenging. Yet, a larger dictionary typically enhances the benefits of WARP-CNN in terms of convergence and solution quality, as demonstrated in Figure 5. Unlike the approach by (Teti et al., 2022), we did not train the backbones on sparse activations but rather opted to reconstruct the image, which allowed us to employ pretrained backbones and evaluate WARP-LCA and LCA as distinct preprocessing modules. Training a classification model on sparse activations obtain by WARP-LCA may provide more robustness.
@@
Looking forward, extending WARP-LCA to directly handle noisy inputs could considerably improve its practical applicability, particularly in scenarios where noise pervades. Additionally, developing adaptive models that can predict not only the initial states but also the optimal number of iterations and lambda may lead to promising results. Such advancements could lead to sparse coding systems that are more autonomous and efficient, effectively bridging the gap between theoretical constructs and their practical implementation.

### v2

@@ 4 Experiments and Results
In our experiments, we used the CIFAR-10 dataset [24], the STL-10 dataset [9], and the Tiny ImageNet dataset [29] for training the models (WARP-CNN). The dictionary was learned using a soft-thresholding (you may find the dictionary in Figure 10). For all of our experiments we normalized the input images to zero-mean and unit variance and also normalized the learned kernels of the dictionary. To train and normalize the dictionary we used the same unsupervised projected gradient based method as in [50]. Additionally, to test the versatility of our method on out-of-distribution data, we employed the Oxford Pets dataset [40] and a high-resolution image from Pexels [37]. For all of our experiments we used the convolutional LCA LCA package  [^1] https://github.com/lanl/lca-pytorch/tree/main presented by Teti et al. [50]. The experiments were conducted on a Linux system equipped with three NVIDIA Tesla V100-PCIE-16GB GPUs, each with a compute capability of 7.0 and 16.94 GB of memory. The system also features 8 CPU cores and memory resources, with a total of 201.18 GB of RAM and 174.20 GB available.
@@
Performance was assessed using Mean Squared Error (MSE), $\ell_{0}$-norm, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM). These metrics were chosen to provide a comprehensive view of both quantitative and qualitative measures. Each model was executed for 1 000 iterations with $\lambda$ set to 0.15 and $\tau$ at 200.
Figure 3: Comparative Speed-Up of WARP-LCA Relative to LCA. The top panels show the maximum PSNR obtained after 1,000 iterations by LCA (red) and WARP-LCA (blue) for CIFAR-10, STL-10, and Tiny ImageNet, illustrating that the PSNR values for both methods typically decrease as the sparsity parameter (i.e., $\lambda$) increases. The bottom panels present the iteration count at which WARP-LCA equals LCA’s PSNR benchmark, alongside the speed-up factor (ratio of LCA’s total iteration count to WARP-LCA’s required count). The speed-up factor generally increases with larger $\lambda$, exhibiting a dataset-dependent minimum within the $\lambda\in[0.15,0.35]$ interval, thereby highlighting WARP-LCA’s accelerated convergence to equivalent reconstruction quality.
To analyze speed improvements, we evaluated models using a $\lambda$ value of 0.15 and a $\tau$ fo 200. Speed comparisons were based on the maximum PSNR achieved by LCA after 1 000 iterations. Since model inference allows control over the number of iterations, we compared both WARP-LCA and standard LCA by evaluating their performance after 1000 iterations to ensure full convergence. The results can be seen in Figure 3, where the WARP method achieved the highest speed up for Tiny Imagenet, where WARP-LCA matches the PSNR of LCA below 100 iterations for all chosen lambdas.
@@
Figure 4: Encoding Performance Comparison between WARP-LCA and LCA at $\lambda=0.15$. The figure displays the performance metrics for WARP-LCA and LCA. The columns represent different metrics: Mean Squared Error (MSE) in log-scale, $\ell_{0}$-Norm, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM). For each plot, WARP-LCA and LCA Mean and Standard Deviation of the Metrics are shown for comparison.
@@
Table 1: Evaluation after last iteration on all datasets with $\lambda=0.15$, $\tau=200$
Metrics LCA Metrics (Mean) WARP-LCA Metrics (Mean)
CIFAR-10
Recon. MSE 0.0112 0.0086
$L_{0}$ Norm 1 274.97 1 232.43
PSNR 30.98 32.15
SSIM 0.95 0.96
STL-10
Recon. MSE 0.0133 0.0095
$L_{0}$ Norm 10 522.21 9 668.79
PSNR 31.68 32.79
SSIM 0.94 0.95
Tiny Imagenet
Recon. MSE 0.0413 0.0328
$L_{0}$ Norm 6 122.32 6 158.49
PSNR 26.82 27.81
SSIM 0.90 0.92
@@
Our results, presented in Figure 4 and in Table 1, showcase the benefits of using WARP-LCA. The WARP-LCA demonstrated markedly faster convergence in terms of MSE, as indicated by the log(MSE) trajectory, maintaining a superior MSE even after 1,000 iterations compared to the standard LCA. Similarly, $\ell_{0}$-norm results showed that WARP-LCA began near optimal sparsity levels, whereas LCA required considerable iterations to approximate these levels. Notably, the $\ell_{0}$-norm achieved by WARP-LCA was lower than that reached by LCA, indicating more efficient sparse coding. Both PSNR and SSIM metrics reflected faster convergence and higher final values for WARP-LCA across all datasets.
This trend was consistent across all datasets, with the only exception for Tiny ImageNet, where the final average sparsity ($\ell_{0}$-norm) was slightly lower for standard LCA.
4.1 Qualitative Comparison of WARP-CNN Outputs and One-Step LCA Refinement
To better understand the qualitative performance of the WARP-CNN, we compared its direct output to the output obtained after a single iteration of the LCA refinement process. This comparison is visualized in Figure 5, where Figure 5(a) displays the direct reconstructions from the WARP-CNN and Figure 5(b) shows the reconstructions after applying one LCA iteration. The datasets evaluated include CIFAR-10, STL-10, and Tiny ImageNet.
The direct output of the WARP-CNN appears notably blurry and exhibits a grayish discoloration across all datasets. Additionally, we observed black artifacts in some reconstructed images, particularly evident in the STL-10 reconstructions. For example, in Figure 5(a), the reconstructions of Image 2 and Image 3 from the STL-10 dataset show visible dark patches and degraded color fidelity.
We hypothesize that these artifacts arise primarily due to the nature of the training cost function and scaling of the targets, which may inadvertently encourage weights that excessively suppress feature activations, resulting in muted reconstructions. With our chosen value for $\gamma=3$ we place a higher emphasis on smaller targets than larger ones.
Interestingly, applying just a single LCA iteration significantly improves the reconstruction quality, as shown in Figure 5(b). The discoloration and artifacts are largely eliminated, and the overall image appears cleaner and more consistent with the original input. Although the reconstruction remains somewhat blurry and lacks certain high-frequency details, the one-step LCA refinement produces images that are noticeably sharper and structurally more accurate. This demonstrates that even minimal iterative refinement can substantially enhance WARP-CNN outputs by correcting initial distortions and reducing artifacts.
(a) WARP-CNN Without LCA. The figure shows images of CIFAR-10, STL-10, and Tiny ImageNet along with the reconstructions obtained from the respective trained WARP-CNN models.
(b) WARP-CNN With One LCA Iteration. The figure shows images of CIFAR-10, STL-10, and Tiny ImageNet and the reconstructions from a single LCA iteration with a state prediction obtained from the respective trained WARP-CNN models. The selected $\tau$ (learning rate) was set to 70.
Figure 5: Side-by-Side Comparison of WARP-CNN Reconstructions. Comparison of WARP-CNN outputs and WARP-CNN with one LCA iteration across CIFAR-10, STL-10, and Tiny ImageNet datasets.
@@ 4.2 Denoising Performance in Classification Pipelines
To evaluate the robustness of our WARP-LCA model in enhancing the denoising capabilities within classification pipelines, we employed three distinct pre-trained convolutional neural network architectures: DenseNet-40 with a growth rate of 12 on CIFAR-10 [20], Wide Residual Network with depth 40 and widening factor 8 on CIFAR-10 [57], and ResNeXt with 29 layers and cardinality 32 using 4D expansions [54].
The test set images from CIFAR-10 were subjected to varying levels of additive Gaussian noise. For each noise level and each backbone architecture, two denoising strategies were applied: traditional LCA-based denoising, run for 800 iterations, and our WARP-LCA-based denoising, run for 200 iterations. This approach allowed us to directly compare the efficacy of our WARP-LCA model against the standard LCA under identical noise conditions, consistent with earlier experiments demonstrating comparable PSNR values between the models. Although different levels of noise typically require adjustments to the sparsity parameter $\lambda$, we opted for a $\lambda$ of 0.2 across all noise levels, drawing from methodologies similar to those reported by Teti et al. [50]. Denoising performance was quantitatively assessed by measuring the classification accuracy post-denoising.
@@
As shown in Figure 6, WARP-LCA consistently outperformed traditional LCA in terms of classification accuracy across all tested noise levels and models.
4.3 Testing WARP-LCA on Larger Images
To evaluate the scalability of WARP-LCA, we selected a smaller model with lower training and validation performance than the original WARP-CNN described in Section 3 (for model details, see the appendix). This choice was driven by GPU memory constraints when processing larger images during evaluation. The test image, a natural scene obtained from Pexels, has dimensions of 1 216×1 824121618241\,216$\times$ 1\,8241 216 × 1 824 pixels.
For this experiment, we used a WARP-LCA model pretrained on the CIFAR-10 dataset, alongside a dictionary similarly trained on CIFAR-10 for comparison with LCA. The results, as depicted in Figure 7 and Figure 8, show a dramatic improvement in convergence speed. All tracked metrics converged even faster than observed on the CIFAR-10, STL10 or Tiny Imagenet test sets.
(a) 100 iterations with $\lambda=0.15$ and $\tau=200$
(b) 100 iterations with $\lambda=0.15$ and $\tau=400$
Figure 7: Comparison of WARP-LCA and LCA reconstructions on a natural scene [37]. The left column shows the original image, the middle image displays the result of WARP-LCA and the right image the result of LCA after 100 iterations with $\lambda=0.15$, $\tau=200$ and 400 respectively.
As shown in 7(a) and the first row of Figure 8 , increasing the learning rate (i.e., decreasing $\tau$) causes standard LCA to converge very slowly and unstably. In contrast, WARP-LCA remains stable and continues to converge rapidly, even with the increased step size. This behavior was consistently observed across experiments on CIFAR-10, STL-10, and Tiny ImageNet, where WARP-LCA effectively handled lower learning rates, while standard LCA using the same dictionary struggled to converge. We attribute this robustness to the high-quality initialization provided by the WARP-CNN, which enables WARP-LCA to tolerate larger step sizes and achieve even faster convergence.
Figure 8: Encoding Performance Comparison between WARP-LCA and LCA at $\lambda=0.15$. The figure displays the performance metrics for WARP-LCA and LCA. The columns represent different metrics: Log Mean Squared Error (MSE), $\ell_{0}$-Norm, Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM). For each plot, WARP-LCA and LCA Mean and Standard Deviation of the Metrics are shown for comparison.
The fully convolutional nature of the trained WARP-CNN facilitates pretraining encoders on smaller datasets and subsequently applying them to much larger images. In such cases, the benefits of using a warm start approach become more pronounced; specifically, the larger the image, the greater the advantage gained from initializing with pretrained encoders. This scalability underscores the effectiveness of WARP-LCA in handling high-resolution, natural images while maintaining efficient convergence behavior.
4.4 WARP-LCA beyond hard thresholding
In this section, we investigate how WARP-LCA behaves when replacing the standard hard thresholding (associated with the $\ell_{0}$ norm) by various alternatives. Specifically, we compare soft thresholding ($\ell_{1}$), hard thresholding ($\ell_{0}$), an $\ell_{1/2}$ approach, and a continuous exact $\ell_{0}$ (CEL0) method on image reconstruction tasks for CIFAR, STL-10, and Tiny ImageNet. The different thresholding and proximal operators were chosen to be the same as in the work of Calatroni et al. [8] (Beyond $\ell_{1}$ Sparse Coding in V1), as their formulations integrate well with the LCA framework. Our primary goal is to analyze reconstruction fidelity and sparsity under each thresholding rule, with metrics including mean squared error (MSE), $\ell_{0}$-norm (approximated by the number of nonzeros), peak signal-to-noise ratio (PSNR), and structural similarity index measure (SSIM). The parameters were set to $\tau=200$ and the algorithm was executed for 100 iterations.
First, the standard hard thresholding (proximal operator for $\ell_{0}$) retains any coefficient $x$ whose magnitude exceeds a threshold $\tau$, while zeroing out smaller coefficients:
$\mathcal{T}_{\text{hard}}(x;\tau)=\begin{cases}x,&\lvert x\rvert\geq\tau,\\[6.0pt] 0,&\lvert x\rvert<\tau.\end{cases}$
Next, the soft thresholding operator ($\ell_{1}$) shrinks coefficients by $\tau$ in magnitude and sets to zero those below the threshold:
$\mathcal{T}_{\text{soft}}(x;\tau)=\mathrm{sign}(x)\,\max\bigl{(}\lvert x\rvert-\tau,\;0\bigr{)}.$
We also employ an $\ell_{1/2}$ thresholding operator as shown in Xu et al. [56], which is more aggressive in promoting sparsity than $\ell_{1}$ proximal operator. A closed-form solution does not appear as straightforward as in the hard or soft cases, yet it can be expressed through the following piecewise definition:
$\Xi_{\theta,1/2}(r^{*})=\begin{cases}f_{\theta,1/2}(r^{*}),&\lvert r^{*}\rvert>\sqrt[3]{54}\,\theta^{2/3},\\[6.0pt] 0,&\text{otherwise},\end{cases}$
where
$f_{\theta,1/2}(r^{*})\;=\;\frac{2\,r^{*}}{3}\,\Bigl{(}1+\cos\Bigl{(}\tfrac{2\pi}{3}\;-\;\tfrac{2}{3}\,\psi_{\theta}(r^{*})\Bigr{)}\Bigr{)},$
and
$\psi_{\theta}(r^{*})\;=\;\arccos\!\Bigl{(}\,\tfrac{\theta}{8}\,\Bigl{(}\tfrac{\lvert r^{*}\rvert}{3}\Bigr{)}^{-\tfrac{3}{2}}\Bigr{)}.$
In practice, the $\ell_{1/2}$ thresholding operator solution as delineated in [31], which presents an algorithm yielding a $k$-sparse solution, is utilized. This approach employs an alternative sparsity criterion compared to selecting a regularization parameter $\lambda$; however, for the purpose of contrasting a “warped” variant with the standard formulation, this methodological difference suffices for comparative analysis.
Finally, we consider the CEL0 approach (Soubies et al. [48]), which serves as a continuous relaxation of $\ell_{0}$ and leads to a thresholding operator that switches between a soft-like behavior and a strict cut-off depending on the norm of the column $\phi_{i}$. Denoting $\lambda$ as a regularization parameter, $\mu$ as an additional scaling factor, and $\|\phi_{i}\|$ as the $\ell_{2}$-norm of the $i$-th dictionary column, the CEL0 threshold can be written as
$\Theta^{\text{CEL0}}_{\mu,\lambda}\bigl{(}r_{i}^{*}\bigr{)}=\begin{cases}\displaystyle\mathrm{sign}(r_{i}^{*})\,\min\Bigl{\{}\,\lvert r_{i}^{*}\rvert,\;\Bigl{(}\lvert r_{i}^{*}\rvert-\sqrt{2\lambda}\,\|\phi_{i}\|\Bigr{)}_{+}\Big{/}\Bigl{(}1\;-\;\tfrac{1}{\|\phi_{i}\|^{2}}\,\mu\Bigr{)}\Bigr{\}},&\|\phi_{i}\|^{2}\,\mu<1,\\[12.0pt] r_{i}^{*}\;\mathbf{1}_{\{\lvert r_{i}^{*}\rvert>\sqrt{2\mu\lambda}\,\|\phi_{i}\|\}}\;+\;\{0,\,r_{i}^{*}\}_{\{\lvert r_{i}^{*}\rvert=\sqrt{2\mu\lambda}\,\|\phi_{i}\|\}},&\|\phi_{i}\|^{2}\,\mu\geq 1.\end{cases}$
Here, $(\,\cdot\,)_{+}$ denotes the positive part, i.e., $(z)_{+}=\max\{0,\,z\}$. This thresholding rule smoothly interpolates between a near-hard thresholding regime and a softening behavior, depending on $\mu$ and the local dictionary norm $\|\phi_{i}\|$.
Having established how each thresholding operator is defined and integrated into the LCA refinement process, we now turn to a thorough experimental comparison on the specified datasets, assessing reconstruction quality and sparsity.
[…]
For CIFAR10, WARP-LCA particularly excels under the $L_{\frac{1}{2}}$ constraint, significantly outperforming LCA by dramatically reducing MSE and substantially improving PSNR and SSIM.
On STL10, WARP-LCA exhibits remarkable superiority under the $L_{0}$ constraint, achieving notably higher PSNR, lower MSE, and enhanced SSIM relative to standard LCA.
[…]
WARP-LCA consistently and significantly improves reconstruction quality across datasets and constraints, with the most pronounced gains observed in the $L_{\frac{1}{2}}$ and $L_{0}$ scenarios. There were almost no instances where standard LCA outperformed WARP-LCA, except for a negligible difference under the $L_{\frac{1}{2}}$ constraint for TinyImg, where LCA showed slightly lower sparsity ($L_{0}$ norm) despite inferior reconstruction quality.
Table 2: Final Metrics Summary Across Datasets and Algorithms
Algorithm MSE $L_{0}$ Norm PSNR SSIM
CIFAR10
$L_{0}$ WARP-LCA 0.0112 1356.51 31.15 0.96
$L_{0}$ LCA 0.0251 1559.46 27.62 0.91
$L_{1}$ WARP-LCA 0.0157 2209.44 29.64 0.94
$L_{1}$ LCA 0.0312 3465.04 26.67 0.90
$L_{\frac{1}{2}}$ WARP-LCA 0.0168 1161.61 29.28 0.93
$L_{\frac{1}{2}}$ LCA 0.2914 2336.29 19.48 0.71
[…]
STL10
$L_{0}$ WARP-LCA 0.0118 10466.21 31.83 0.94
$L_{0}$ LCA 0.0923 12281.02 24.91 0.83
$L_{1}$ WARP-LCA 0.0152 18532.00 30.71 0.92
$L_{1}$ LCA 0.0277 31702.58 28.07 0.88
$L_{\frac{1}{2}}$ WARP-LCA 0.1812 1713.08 20.00 0.62
$L_{\frac{1}{2}}$ LCA 0.5954 1633.49 15.19 0.43
[…]
TinyImg
$L_{0}$ WARP-LCA 0.0398 6680.87 27.07 0.90
$L_{0}$ LCA 0.0992 6686.10 23.45 0.80
$L_{1}$ WARP-LCA 0.0491 10006.09 26.13 0.88
$L_{1}$ LCA 0.0839 15048.11 23.79 0.80
$L_{\frac{1}{2}}$ WARP-LCA 0.1833 861.47 19.89 0.61
$L_{\frac{1}{2}}$ LCA 0.4952 808.62 16.08 0.46
[…]
@@ 5 Discussion
The comparative analysis in Section 4 (illustrated in Figure 4) compellingly demonstrates the superior efficacy of WARP-LCA, over the traditional LCA: faster convergence and settling for deeper minima. This attribute was consistently evident across a range of noise levels and architectural backbones in the denoising experiment in Section 4.1. These results not only corroborate but also extend the findings by Teti et al. [50], propelling WARP-LCA beyond the current state-of-the-art. This advancement underscores the potential of predictive initializations to redefine performance benchmarks in the field, offering robust, fast, and more accurate sparse coding.
The robustness of WARP-LCA, particularly for out-of-distribution data, was tested on the Oxford Pets dataset [40], which starkly contrasts with CIFAR-10 in terms of image size (images in the Oxford Pets datset are much larger) and image statistics. The Oxford Pets dataset is a collection of images featuring 37 different breeds of cats and dogs, each annotated with class labels, bounding boxes, and pixel-level segmentation masks. Despite these considerable differences, and the fact that our models were initially trained only on CIFAR images, the fully convolutional nature of predictor architectures allowed for an efficient adaptation to this new context.
@@
The (relative) accumulated activation map generated by WARP-LCA provides more homogeneous and less extreme structures than the accumulated LCA activation map, indicated much more uniformly distributed activations. Even after 300 iterations, activity in the accumulated LCA map is still concentrated in very dark and light areas of the image. In addition, WARP-LCA also has a lower $\ell_{0}$-Norm than the LCA. The reconstructed image from the LCA appears not to have converged yet, as the colors, especially in the light and dark areas, seems to be faded. The reconstructed image from the WARP-LCA seems to be much closer to the original.
This may imply that the CIFAR dictionary alone may be sufficient to be used on data other than CIFAR images and WARP-LCA managed to predict states that lead to a consistent and balanced activation profile.
We have experimented with different approaches to warm starting LCA, including random initialization but found that the only methods that lead to convergence are the conventional zero initialization as is typical for LCA and our CNN predictor. We have compared the similarity of learned cifar10-kernels to the correlation of activations of LCA and WARP-LCA after 1000 optimization cycles up to convergence and found that the WARP-LCA method has lower correlation between kernels as would be expected from the earlier results. The prediction of initial states leads to more diverse and robust minima (please see the appendix for a comparrison of LCA, WARP-LCA, CNN-only prediction and simple convolution).
@@
It is important to note that our experiments were conducted using a single, learned over-complete dictionary. The application of WARP-LCA to dictionaries with lower overcompleteness—and hence fewer kernels—might prove more challenging. Yet, a larger dictionary typically enhances the benefits of WARP-CNN in terms of convergence and solution quality, as demonstrated in Figure 9. Unlike the approach by [50], we did not train the backbones on sparse activations but rather opted to reconstruct the image, which allowed us to employ pretrained backbones and evaluate WARP-LCA and LCA as distinct preprocessing modules. Training a classification model on sparse activations obtain by WARP-LCA may provide more robustness.
@@
5.1 Limitations
While our results underscore the potential of WARP-LCA, several limitations remain. In WARP-LCA, a convolutional neural network (CNN) first provides a warm start for the sparse coefficients, after which an iterative LCA procedure refines those coefficients. This two-stage process naturally incurs additional computational cost relative to running LCA alone or a standalone CNN. Specifically, the input must be passed through the CNN once, and then multiple LCA iterations are performed, each updating the sparse coefficients. The computational overhead becomes more pronounced as input sizes grow or the number of refinement steps increases.
As a concrete example, in Section 4.3 we had to reduce the size of the trained CNN for larger images so that it would fit in GPU memory; using the full architecture would have exceeded our hardware resources. We did not exhaustively explore potential network architectures or perform a broad hyperparameter grid search to identify more parameter-efficient (or memory-friendly) configurations. Consequently, the scalability of WARP-LCA—especially in terms of network depth, input resolution, and number of iterative updates warrants deeper investigation.
In addition, future efforts could extend the WARP-LCA framework to deep or hierarchical sparse coding models [7], including deep convolutional variants, thereby enabling richer feature hierarchies and more expressive representations. We also have not explored the full range of contemporary LCA applications, such as robust model inversion attacks [12], on-chip learning [49], or robust audio classification [14]. Given that WARP-LCA demonstrates improvements across multiple algorithms and datasets, it is very likely these advantages would also translate effectively into such domains. Lastly, further work on adaptive models that not only estimate the initial states but also automatically learn hyperparameters—such as the optimal number of iterations and the sparsity penalty parameter $\lambda$—could help reduce manual tuning, improving both efficiency and practicality. These advancements would help bridge the gap between the current theoretical framework and large-scale deployment scenarios.